OpenStudy (anonymous):

Solve the equation, if possible. (If there is no solution, enter NO SOLUTION.) 8x + 8 = 7x + 3 x= an identity an equation with no solutions a conditional equation

OpenStudy (anonymous):

x = -5

jimthompson5910 (jim_thompson5910):

8x + 8 = 7x + 3 8x = 7x + 3-8 8x - 7x = 3-8 x = -5 So the answer is choice C

OpenStudy (anonymous):

conditional equation?

jimthompson5910 (jim_thompson5910):

yes this is a conditional equation